OpenStudy (anonymous):

Simplify. The square root of 27/12.

OpenStudy (anonymous):

okkk so simplify \[ \sqrt{ \frac{27}{12}}\]

OpenStudy (anonymous):

SHamil gtfo out of here

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

O_O? lol

OpenStudy (shamil98):

\[\sqrt{\frac{ 27 }{ 12 }} = \sqrt{\frac{ 9*3 }{ 4*3 }} = ?\] simplify it. there are two numbers that you can take out of the square root symbol and two that will stay inside the symbol.

OpenStudy (anonymous):

Oh, okay.

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ 27\implies 3\cdot 3\cdot 3\implies 3^2\cdot 3\qquad 12\implies 2\cdot 2\cdot 3\implies 2^2\cdot 3}}\\ \quad \\ \quad \\ \sqrt{\cfrac{27}{12}}\implies \cfrac{\sqrt{27}}{\sqrt{12}}\implies \cfrac{\sqrt{3^2\cdot 3}}{\sqrt{2^2\cdot 3}}\)

OpenStudy (anonymous):

So, the 9 and the 4 stay in, right?

OpenStudy (jdoe0001):

\(\bf 9 = 3^2 \quad 4 = 2^2\)

OpenStudy (anonymous):

Oh! Thank you! :)

OpenStudy (shamil98):

Simplifying it out what would your final answer be?

OpenStudy (anonymous):

Sorry, the images wouldn't come up. It would be 3 over 2 right?

OpenStudy (anonymous):

ok first break apart the square root. square root of 27=3root3 square root of 12=2root3